OpenStudy (anonymous):

The perimeter of a triangle is 199 inches. If one side of the triangle is six more than the shorter side, and the longer side is 19 more than the shorter side, find the lengths of the three sides?

OpenStudy (anonymous):

short side = x one side = x + 6 longer side = x + 19

OpenStudy (anonymous):

Perimeter = Side + side + side

OpenStudy (anonymous):

199= (x) + (x+6) + (x+19)

OpenStudy (anonymous):

@huntieee

OpenStudy (anonymous):

Understand so far @huntieee

OpenStudy (anonymous):

58,64,77